40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 13 14:26:00 2022, JC, Update, Alignment, LO2 Beam 
    Reply  Fri May 13 18:42:13 2022, Paco, Update, Alignment, ITMX and ITMY sat amp failures 
Message ID: 16852     Entry time: Fri May 13 18:42:13 2022     In reply to: 16851
Author: Paco 
Type: Update 
Category: Alignment 
Subject: ITMX and ITMY sat amp failures 

[Yuta, Anchal, Paco]

As described briefly by JC, there were multiple failure modes going during this work segment. blush


ITMX SatAmp SAGA

Indeed, the 64 pin crimp cable from the gold sat amp box broke when work around ITMX chamber was ongoing. We found the right 64 pin head replacement around and moved on to fix the connector in-situ. After a first attempt, we suddenly lost all damping on vertex SUS (driven by these old sat amp electronics) because our c1susaux acromag chassis stopped working. After looking around the 1x5 rack electronics we noted that one of the +- 20 VDC Sorensens were at 11.6 VDC, drawing 6.7 A of current (nominally this supply draws over 5 Amps!) so we realized we had not connected the ITMX sat amp correctly, and the DC rail voltage drop busted the acromag power as well, tripping all the other watchdogs devil ...

We fixed this by first, unplugging the shorted cable from the rack (at which point the supply went back to 20 VDC, 4.7 A) and then carefully redoing the crimp connector. The second attempt was successful and we restored the c1susaux modbusIOC service (i.e. slow controls).


ITMY SatAmp SAGA

As we restored the slow controls, and damped most vertex suspensions, we noticed ITMY UL and SD osems were reading 0 counts both on the slow and fast ADCs. crying We suspected we had pulled some wires around when busy with the ITMX sat amp saga. We found that Side OSEM cLEMO cable was very loose on the whitening board. In fact, we have had no side osem signal on ITMY for some time. We fixed this. Nevertheless the UL channel remained silent... We then did the following tests:

  • Test PD mon outputs on the whitening card. We realized the whitening cards were mislabeled, with ITMX and ITMY flipped angry. We have labeled them appropriately.
  • Tested input DB15 cable with breakout board.
  • Went to the ITMY sat amp box and used the satellite box TESTER 2 on J1. It seemed correct.
  • We opened the chamber, tested the in-vacuum segments, they all were ok.
  • We flipped UR-UL OSEMs and found that the UL OSEM is healthy and works fine on UR channel.
  • We tested the in-air cable between satellite box and vacuum flange and it was ok too.
  • We suspected that the satellite box tester  is lying, so we replaced the satellite box with the spare old MC1 satellite box, and indeed that solved the issue.

DO NOT TRUST THE SATELLITE BOX TESTER 2.


Current state:

  • IMC locking normally.
  • All suspensions are damping properly.
  • Oplevs are not centered.
  • No flashing on either of the arms. We had no luck in ~20 min of attempt with just input injection changed.
  • On kicking PR3, we do see some flashing on XARM, which means XARM cavity atleast is somewhat aligned.
  • All remaining tasks before pumpdown are still remaining. We just lost the whole day.
ELOG V3.1.3-